3. Jarett & Sons's common stock currently trades at $28.00 a share. It is expected to...

3.

Jarett & Sons's common stock currently trades at $28.00 a share. It is expected to pay an annual dividend of $2.25 a share at the end of the year (D1 = $2.25), and the constant growth rate is 7% a year.

  1. What is the company's cost of common equity if all of its equity comes from retained earnings? Do not round intermediate calculations. Round your answer to two decimal places.

      %

  2. If the company issued new stock, it would incur a 15% flotation cost. What would be the cost of equity from new stock? Do not round intermediate calculations. Round your answer to two decimal places.

      %
Answer #1

a)

Cost of equity = (D1 / price) + growth

Cost of equity = (2.25 / 28) + 0.07

Cost of equity = 0.080357 + 0.07

Cost of equity = 0.1504 or 15.04%

2)

Flotation cost = 15% of 28 = 4.2

Price of flotation cost = 28 - 4.2 = 23.8

Cost of new equity = (2.25 / 23.8) + 0.07

Cost of new equity = 0.094538 + 0.07

Cost of new equity = 0.1645 or 16.45%